Suppose a​ monopoly's price is

​$60.00

and its marginal cost of production is

​$48.00

What is the​ firm's markup

​?

The​ monopoly's markup is

percent. ​(Enter a numeric response using a real number rounded to two decimal​places.)
Answer #1

Firms markup= P-MC = 60-48 = 12

The monopoly's Markup is (P-MC)*100/MC = 12*100/48 = 25